Circular No. 6932

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1748–288 D. A. Smith, A. Levine, and A. Wood, Massachusetts Institute of Tech-

nology (MIT), report for the RXTE ASM team at MIT and NASA Goddard Space Flight Center: “The RXTE All-Sky Monitor has detected a transient x-ray source at α = 17^{h}48^{m}08^{s}, δ = −28^{o}28′.8 (equinox 2000.0; error radius 5′, to better than 90-percent confidence). The source was observed in each of two detector systems in two contiguous 90-s exposures on June 4.69 UT, and again in two more observations 1.8 hr later. The spectrum is hard, with typical fluxes of 90, 300, and 470 mCrab in the bands 1.5–3, 3–5, and 5–12 keV, respectively.”

XTE J2012+381 R. I. Hynes and P. Roche, University of Sussex, report: “Images were

taken with the 1.0-m Jacobus Kapteyn Telescope, La Palma, on June 3.14– 3.23 UT by N. Walton (Isaac Newton Group) of USNO 1275.13846761 (A1.0 catalogue position α = 20^{h}12^{m}37^{s}.80, δ = +38^{o}11′00′′.6, equinox 2000.0), a possible optical counterpart (IAUC 6931) to XTE J2012+381 (IAUC 6920). We estimate U = 19.4, B = 19.1, V = 17.9, R = 17.2, I = 16.6 (± 0.2 mag). V, R, and I images also show a faint red companion star 1′′.2 away; we measure an offset in α of −0^{s}.09 ± 0^{s}.01, and in δ of +0′′.5 ± 0′′.1 (i.e., position end figures 37^{s}.71, 01′′.1). We estimate V = 21.3, R = 20.1, I = 18.6 (± 0.3 mag). These colors are consistent with a reddened soft x-ray transient, as suggested by the estimate of A(V) = 7 deduced from ASCA spectral fitting (IAUC 6927). We therefore propose that this object is the true optical counterpart to XTE J2012+381.”

R. M. Hjellming and M. P. Rupen, National Radio Astronomy Observatory (NRAO); and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “Very Large Array observations on June 4.38 UT of the radio candidate for the x-ray transient XTE J2012+381 (cf. IAUC 6924) showed flux densities of 3.5, 3.0, and 2.6 mJy at 1.4, 4.9, and 8.4 GHz, respectively. Since the 1.4- and 4.9-GHz data are now significantly greater than on May 31.25 UT (IAUC 6924), and because of the variability at 15 GHz reported by Pooley (IAUC 6926), this source is very likely the radio counterpart of XTE J2012+381. We can also report an improved position of α = 20^{h}12^{m}37^{s}.67, δ = +38^{o}11′01′′.2 (equinox 2000.0; uncertainty 0′′.4).”

1998 June 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green